OpenStudy (anonymous):

19.a bird flies using v=| t-2| find the distance and displacement covered by bird in 4 second with graph?

OpenStudy (akashdeepdeb):

velocity is given by = |t-2| We have to find the distance and displacement in 4 seconds. [t = 4] Thus v = |t-2| = |4-2| = 2m/s t = 4s x = vt = 8m [Displacement] You may find the distance by drawing the graph. Take t along the x-axis v which is a function of time along the y-axis  This, must be the answer. @LastDayWork Is this correct?